# Multi-Layer Transformers Gradient Can be Approximated in Almost Linear Time

### v1

@@ 3.3 Close forms of gradient components
Then, we are ready to introduce the close forms of the three gradient components in a single self-attention transformer layer. Notably, according to the chain rule, the gradient of the $k$-th transformer layer in LLMs depends on the gradient components from the $(k+1)$-th transformer layer. The gradient can be calculated for every transformer layer by combining the upstream and local gradients. The close forms of the gradients for each layer in multi-layer transformers are formalized in the following lemma (Lemma 3.4).
Lemma 3.4 (Close form of gradient components, informal version of Lemma C.3).
Let $L(X)$ be defined as Definition 3.1. Let $W_{i}=W_{Q_{i}}W_{K_{i}}^{\top},W_{V_{i}}\in\mathbb{R}^{d\times d}$ denote the attention weight in the $i$-th transformer layer. Let $T_{i}(X)$ denote the intermediate variable output by $i$-th self-attention transformer layer (see Definition 3.3). Let $G_{i}\in\mathbb{R}^{n\times d}$ denote the gradient matrix resulting from the application of the chain rule up to the function $g_{i}$, i.e., $G_{i}=\frac{\mathrm{d}L(X)}{\mathrm{d}\mathsf{Attn}_{i}(T_{i-1}(X))}$. For $j\in[n],k\in[d]$, let $G_{i}(j,k)$ denote the $(j,k)$-th entry of $G_{i}$, let $\frac{\mathrm{d}\mathsf{Attn}_{i}(T_{i-1}(X))_{j,k}}{\mathrm{d}T_{i-1}(X)}\in\mathbb{R}^{n\times d}$ denote the gradient of $(j,k)$-th entry of $\mathsf{Attn}_{i}(T_{i-1}(X))$. Then, we can show that
@@
  • Part 2.
    $\frac{\mathrm{d}L(X)}{\mathrm{d}W_{i}}=\sum_{j=1}^{n}\sum_{k=1}^{d}G_{i}(j,k)\cdot\frac{\mathrm{d}\mathsf{Attn}_{i}(T_{i-1}(X))_{j,k}}{\mathrm{d}W_{i}}.$
  • Part 3.
    $\frac{\mathrm{d}L(X)}{\mathrm{d}W_{V_{i}}}=\sum_{j=1}^{n}\sum_{k=1}^{d}G_{i}(j,k)\cdot\frac{\mathrm{d}\mathsf{Attn}_{i}(T_{i-1}(X))_{j,k}}{\mathrm{d}W_{V_{i}}}.$
Based on the above close forms of three gradient components, we can move on to our main results smoothly.
4 Main Result
@@ 4.1 Fast computing for single layer
We assume $d=O(\log n)$ and each number in matrices can be written using $O(\log n)$ bits. Let $L(X)$ be defined as Definition 3.1. Suppose we have a single-layer self-attention transformer model ($m=1$ in Definition 1.3). We can approximate one-layer self-attention for three gradient components, i.e. $\frac{\mathrm{d}L(X)}{\mathrm{d}X}$, $\frac{\mathrm{d}L(X)}{\mathrm{d}W}$ and $\frac{\mathrm{d}L(X)}{\mathrm{d}W_{V}}$, in $n^{1+o(1)}$ time with $1/\operatorname{poly}(n)$ error.
@@
Algorithm 1 Almost Linear Time (ALT) Multi-layer Transformer Gradient Approximation Algorithm
@@
9:     $W_{i},W_{V_{i}}\in\mathbb{R}^{d\times d}$ : the weight matrices in $i$-th transformer layer
@@
17:      /* Approximate $\frac{\mathrm{d}L(X)}{\mathrm{d}W_{i}}$, Lemma 5.2 */
@@
23:     return $\widetilde{g}_{t},\widetilde{g}_{w},\widetilde{g}_{v}$triangleright $\widetilde{g}_{t},\widetilde{g}_{w},\widetilde{g}_{v}$ are the approximated $\frac{\mathrm{d}L(X)}{\mathrm{d}T_{i-1}},\frac{\mathrm{d}L(X)}{\mathrm{d}W_{i}},\frac{\mathrm{d}L(X)}{\mathrm{d}W_{V_{i}}}$ in next iteration, respectively.
@@
31:         Optimize $W_{i}$ via $\widetilde{g}_{w}$ using optimizer
@@ 4.2 Fast computing for multi-layer transformers
Based on the results demonstrated in previous sections, we are ready to introduce our main result: the whole transformer model can be approximated in almost linear time.
@@
Let $m$ denote the number of transformer layers. We assume $d=O(\log n)$ and each number in matrices can be written using $O(\log n)$ bits. We can show that, for any $i\in[m]$, all the gradient components (see also Lemma 3.4) of the $i$-th layer can be computed by Algorithm 1 in almost linear time $n^{1+o(1)}$, and the approximation error of entire $m$ layer transformer model can be bounded by $1/\operatorname{poly}(n)$.
@@
The theorem can be proved by directly combining Theorem 4.1 and Lemma 5.5. ∎
Theorem 4.2 demonstrates that, during the training of a multi-layer transformer model, at each training iteration, the gradient computation for the weight matrices of each layer can be performed in almost linear time $n^{1+o(1)}$. This result supports the feasibility of fast training for any transformer-based large language models (LLMs). In Algorithm 1, we illustrate the process of back-propagating gradients from the $m$-th transformer layer back to the first layer. This algorithm highlights the significance of the gradient with respect to the intermediate variables $T_{i}(X)$. Due to the application of the chain rule in gradient computation, the gradient of $T_{i}(X)$ is indispensable for determining the gradients of the weight matrices $W_{i}$ and $W_{V_{i}}$ at the $i$-th layer. Consequently, by iteratively computing the gradient for $T_{i}(X)$, we systematically propagate the gradient through to the initial transformer layer.
Additionally, our algorithm is capable of computing the gradient with respect to the input data $X$. Therefore, our algorithm also supports fast prompt tuning. For a more in-depth discussion on this topic, please refer to Section 6.
4.3 Beyond the previous works
Our algorithm exhibits significant advancements over two brilliant related prior studies, namely [5] and [6]. In [5], the authors proposed an almost linear time algorithm for computing the forward process of the attention mechanism. In contrast, [6] introduced an almost linear time algorithm for the backward of attention mechanism. However, [6] has the following limitations: (1) only computing gradients for a single layer of the attention mechanism, which cannot extend to multiple layers; (2) calculating gradients with respect to a specific loss, namely the $\ell_{2}$ loss; (3) computing gradients only for the weight matrix $W_{i}$ (as defined in Definition 1.2), but ignore other crucial components such as the MLP layer following attention computation and the activation function.
In our work, we have the following improvements beyond [6]: (1) we enable almost linear time gradient computation across an entire transformer layer, incorporating both the MLP layer and the activation function; (2) our algorithm supports gradient calculation for any loss function $L(X)$ (see Definition 3.1); (3) we extend the gradient calculation to include not only $W_{i}$ but also $T_{i}(X)$ and $W_{V_{i}}$. These advancements collectively demonstrate a substantial leap forward from the methodologies in [5] and [6].
@@ 5 Technical Overview
This section provides a brief overview of the proof techniques used throughout this paper. Section 5.1 delves into intuition under the low-rank approximation applied to the attention matrix $f(X)$. Section 5.2 elaborates on calculating gradients with respect to the intermediate variable (hidden states) $T_{i}$ in almost linear time. In Section 5.3, we detail the process for computing gradients for the weight matrices $W$ and $W_{V}$ within transformer layers in almost linear time. Finally, in Section 5.4, we demonstrate how to generalize our findings from single-layer transformers to multi-layer transformers, including an analysis of the runtime and approximation error across the entire model.
@@ 5.1 Low-rank approximation for attention matrix
In this section, we delve into the crucial technique behind our work: the low-rank approximation of the attention matrix, which is achieved through the polynomial method [3, 1]. Drawing inspiration from [5], the intuition of this approximation lies in the fact that the attention matrix $f(X)\in\mathbb{R}^{n\times n}$ (as defined in Definition 1.2), also referred to as the similarity matrix in attention mechanism, can be effectively approximated by low-rank matrices $U_{1},V_{1}\in\mathbb{R}^{n\times k_{1}}$, where $k_{1}=n^{o(1)}$. The naive method for calculating the attention matrix $f(X)$ has a time complexity of $O(n^{2})$, whereas the input data $X\in\mathbb{R}^{n\times d}$ contains only $d\cdot n=n^{1+o(1)}$ entries. This discrepancy suggests the potential of using low-rank representations of $f(X)$ to design a fast algorithm.
An example of how to use the low-rank representations is the attention forward ($\mathsf{Attn}(X):=f(X)V$ in Definition 1.2) as in [5]: approximating $f(X)$ along does not lead to fast algorithm, since $U_{1}V_{1}^{\top}$ still requires $n\times n$ entries. But by using the structure of the whole, we can do it faster. By expressing $f(X)$ as $U_{1}V_{1}^{\top}$, the attention forward becomes $\underbrace{U_{1}}_{n\times k_{1}}\underbrace{V_{1}^{\top}}_{k_{1}\times n}\underbrace{V}_{n\times d}.$ It is well known that different multiplication sequences can lead to dramatically different numbers of operations required, so the order of matrix multiplications matters. We first perform $V_{1}^{\top}V\in\mathbb{R}^{k_{1}\times d}$ and this cost $O(k_{1}nd)=n^{1+o(1)}$ time. Then we perform $U_{1}V_{1}^{\top}V$ costing $O(nk_{1}d)=n^{1+o(1)}$ time.
This method significantly reduces the computation time of the attention forward from $O(n^{2})$ to almost linear time, $n^{1+o(1)}$. Driven by this technique and analyzing the close forms of the gradients, we can extend the acceleration to the gradient of the entire model.
@@ 5.2 Accelerating gradient computation of $T_{i}(X)$
Based on the low-rank approximation method mentioned in Section 5.1, we can compute the gradient of $L(X)$ with respect to the intermediate variable $T_{i}(X)$, which denotes the output of the $i$-th transformer layer. This computation is critical as it enables us to calculate gradients for other gradient components because of the chain rule.
Extending to any kind of loss function.
According to the findings in [25], the gradient $\frac{\mathrm{d}L(X)}{\mathrm{d}T_{i}(X)}$ can be decomposed into five components, namely $C_{6}(X),C_{7}(X),C_{8}(X),C_{2}(X),C_{4}(X)$, as detailed in Lemma D.1. However, the gradient result presented in [25] is tailored to a specific loss function, the $\ell_{2}$ loss, limiting its applicability to a narrow range of scenarios. In this study, we extend the scope of their findings by extending them to apply to any loss function $L(X)$, as defined in Definition 3.1. By incorporating $G_{i}=\frac{\mathrm{d}L(X)}{\mathrm{d}\mathsf{Attn}_{i}(T_{i-1}(X))}$, we derive a closed-form expression for the gradient of $L(X)$ with respect to $T_{i}(X)$, which is detailed in Section D.2.
@@
To accelerate the gradient computation for $T_{i}(X)$, we need the matrix form of the gradients, as discussed in Section D. This approach is essential for understanding the underlying mechanisms when applying the low-rank approximation technique in gradient calculations. Subsequently, using that technique, we can accelerate the gradient computation for $\frac{\mathrm{d}L(X)}{\mathrm{d}T_{i}(X)}$ (Lemma 5.1). By individually applying this technique to each of the five terms, we demonstrate that each term can be computed in almost linear time, $n^{1+o(1)}$, as shown in Sections E.1, E.2, E.3, E.4, and E.5.
The next step is to aggregate these terms, as described in Section E.6. Since all five terms are $n\times d$ matrices, the summation of these terms remains almost linear in complexity. Conclusively, we are safe to argue that for any single-layer transformer, the gradient computation with respect to the input tensor can be performed in almost linear time $n^{1+o(1)}$, as stated in Lemma 5.1.
The statement made for a single transformer layer can be readily generalized to any layer within an $m$-layer transformer model. For instance, consider the intermediate variables $T_{i}(X)$ and $T_{i-1}(X)$ (as defined in Definition 3.3), where $T_{i}(X)=(g_{i}\circ\mathsf{Attn}_{i})(T_{i-1}(X))$. Given the gradient $\frac{\mathrm{d}L(X)}{\mathrm{d}T_{i}(X)}$, as established in the previous paragraph, we can compute the gradient with respect to $T_{i-1}(X)$, namely $\frac{\mathrm{d}L(X)}{\mathrm{d}T_{i-1}(X)}$, in almost linear time $n^{1+o(1)}$. For a multi-layer transformer model, the above process can be conducted recursively. Thus, we can compute the gradient of the loss function $L(X)$ on any $T_{i}(X)$ in almost linear time $n^{1+o(1)}$.

### v2

@@ 3.2 Closed forms of gradient components
Then, we are ready to introduce the closed forms of the three gradient components in a single self-attention transformer layer. Notably, according to the chain rule, the gradient of the $k$-th transformer layer in LLMs depends on the gradient components from the $(k+1)$-th transformer layer. The gradient can be calculated for every transformer layer by combining the upstream and local gradients. The closed forms of the gradients for each layer in multi-layer transformers are formalized in the following lemma (Lemma 3.4).
Lemma 3.4 (Closed form of gradient components, informal version of Lemma C.4).
Let $L(X)$ be defined as in Definition 3.1, and the $m$-layer transformer defined as in Definition 1.3. Let $W_{Q_{i}},W_{K_{i}},W_{V_{i}}\in\mathbb{R}^{d\times d}$ denote the attention weight in the $i$-th attention. Let $T_{i}(X)$ denote the intermediate variable output by $i$-th self-attention transformer layer (see Definition 3.3). Let $G_{i}\in\mathbb{R}^{n\times d}$ denote the gradient matrix resulting from the application of the chain rule up to the function $g_{i}$, i.e., $G_{i}=\frac{\mathrm{d}L(X)}{\mathrm{d}\mathsf{Attn}_{i}(T_{i-1}(X))}$. For $j\in[n],k\in[d]$, let $G_{i}(j,k)$ denote the $(j,k)$-th entry of $G_{i}$, let $\frac{\mathrm{d}\mathsf{Attn}_{i}(T_{i-1}(X))_{j,k}}{\mathrm{d}T_{i-1}(X)}\in\mathbb{R}^{n\times d}$ denote the gradient of $(j,k)$-th entry of $\mathsf{Attn}_{i}(T_{i-1}(X))$. Then, we can show that
@@
  • Part 2. Let $W_{*_{i}}$ be $W_{Q_{i}},W_{K_{i}}$ or $W_{V_{i}}$, then
    $\frac{\mathrm{d}L(X)}{\mathrm{d}W_{*_{i}}}=\sum_{j=1}^{n}\sum_{k=1}^{d}G_{i}(j,k)\cdot\frac{\mathrm{d}\mathsf{Attn}_{i}(T_{i-1}(X))_{j,k}}{\mathrm{d}W_{*_{i}}}.$
Our main results are based on the above closed forms of four gradient components.
4 Main Results
@@ 4.1 Fast computing for single layer
We assume $d=O(\log n)$ and each number in matrices can be written using $O(\log n)$ bits. Assume the number of layers $m=n^{o(1)}$. Let $L(X)$ be defined as Definition 3.1. Suppose we have a single-layer self-attention transformer model ($m=1$ in Definition 1.3). We can approximate one-layer self-attention for three gradient components, i.e. $\frac{\mathrm{d}L(X)}{\mathrm{d}X}$, $\frac{\mathrm{d}L(X)}{\mathrm{d}W_{Q}W_{K}^{\top}}$ and $\frac{\mathrm{d}L(X)}{\mathrm{d}W_{V}}$, in $n^{1+o(1)}$ time with $1/\operatorname{poly}(n)$ error.
@@
Algorithm 1 Almost Linear Time (ALT) Multi-layer Transformer Gradient Approximation
@@
9:     $W_{Q_{i}},W_{K_{i}},W_{V_{i}}\in\mathbb{R}^{d\times d}$ : the weight matrices in $i$-th transformer layer
@@
17:      /* Approximate $\frac{\mathrm{d}L(X)}{\mathrm{d}W_{Q_{i}}W_{K_{i}}^{\top}}$, Lemma 5.2 */
@@
23:     return $\widetilde{g}_{t},\widetilde{g}_{w},\widetilde{g}_{v}$triangleright $\widetilde{g}_{t}$ is the approximated $\frac{\mathrm{d}L(X)}{\mathrm{d}T_{i-1}}$ for back-propagation
@@
31:         Optimize $W_{Q_{i}},W_{K_{i}}$ via $\widetilde{g}_{w}$ using optimizer
@@ 4.2 Fast computing for multi-layer transformers
Based on the results demonstrated in previous sections, we are ready to introduce our main result: the gradients of the whole transformer model can be approximated in almost linear time.
@@
Let $m$ denote the number of transformer layers. We assume $d=O(\log n)$ and each number in matrices can be written using $O(\log n)$ bits. Assume the number of layers $m=n^{o(1)}$. We can show that, for any $i\in[m]$, all the gradient components (see also Lemma 3.4) of the $i$-th layer can be computed by Algorithm 1 in almost linear time $n^{1+o(1)}$, and the approximation error of the entire $m$ layer transformer model can be bounded by $1/\operatorname{poly}(n)$.
@@
We prove the theorem by directly combining Theorem 4.1 and Lemma 5.5. ∎
Theorem 4.2 demonstrates that, during the training of a multi-layer transformer model, at each training iteration, the gradient computation for the weight matrices of each layer can be performed in almost linear time $n^{1+o(1)}$. This result supports the feasibility of fast training for any transformer-based large language models (LLMs). In Algorithm 1, we illustrate the process of back-propagating gradients from the $m$-th transformer layer back to the first layer. This algorithm highlights the significance of the gradient with respect to the intermediate variables $T_{i}(X)$. Due to the application of the chain rule in gradient computation, the gradient of $T_{i}(X)$ is indispensable for determining the gradients of the weight matrices $W_{Q_{i}},W_{K_{i}}$ and $W_{V_{i}}$ at the $i$-th layer. Consequently, by iteratively computing the gradient for $T_{i}(X)$, we systematically propagate the gradient through to the initial transformer layer. Additionally, our algorithm is capable of computing the gradient with respect to the input data $X$. Therefore, our algorithm also supports fast prompt tuning. For a more in-depth discussion on this topic, please refer to Section 6.
4.3 Beyond the previous work
Our algorithm exhibits significant advancements over two seminal prior studies, namely AS [5] and AS24a [6]. In AS [5], the authors proposed an almost linear time algorithm for computing the forward process of the attention mechanism. In contrast, AS24a [6] introduced an almost linear time algorithm for the backward of attention mechanism. However, AS24a [6] has the following limitations: (1) only computing gradients for a single layer of the attention mechanism, which cannot extend to multiple layers; (2) calculating gradients with respect to a specific loss, namely the $\ell_{2}$ loss; (3) computing gradients only for the weight matrix $W_{Q_{i}},W_{K_{i}}$ (as defined in Definition 1.2), but ignore other crucial components such as the MLP layer following attention computation and the activation function. These limitations are inherent in their technique and prevents the applicability of the method in multiple layer transformers.
In our work, we have the following improvements beyond previous work: (1) we enable almost linear time gradient computation across an entire transformer layer, incorporating both the MLP layer and the activation function; (2) our algorithm supports gradient calculation for general loss function $L(X)$ (see Definition 3.1); (3) we extend the gradient calculation to include not only $W_{Q_{i}},W_{K_{i}}$ but also $T_{i}(X)$ and $W_{V_{i}}$. These advancements collectively demonstrate a substantial leap forward from the methodologies in AS [5] and AS24a [6].
@@ 5.1 Low-rank approximation for attention matrix
In this section, we delve into the crucial techniques behind our work: the low-rank approximation of the attention matrix, which is achieved through the polynomial method [3, 1]. Drawing inspiration from AS [5], the intuition of this approximation lies in the fact that the attention matrix $f(X)\in\mathbb{R}^{n\times n}$ (as defined in Definition 1.2), also referred to as the similarity matrix in attention mechanism, can be effectively approximated by low-rank matrices $U_{1},V_{1}\in\mathbb{R}^{n\times k_{1}}$, where $k_{1}=n^{o(1)}$. The naive method for calculating the attention matrix $f(X)$ has a time complexity of $O(n^{2})$, whereas the input data $X\in\mathbb{R}^{n\times d}$ contains only $d\cdot n=n^{1+o(1)}$ entries. This discrepancy suggests the potential of using low-rank representations of $f(X)$ to design a fast algorithm.
An example of how to use the low-rank representations is the attention forward. First note that approximating $f(X)$ alone does not lead to a fast algorithm, since $U_{1}V_{1}^{\top}$ still requires $n\times n$ entries. But by using the structure of the attention $\mathsf{Attn}(X):=f(X)V$ where $V=XW_{V}$, we can do it faster. By expressing $f(X)$ as $U_{1}V_{1}^{\top}$, the attention forward becomes $\underbrace{U_{1}}_{n\times k_{1}}\underbrace{V_{1}^{\top}}_{k_{1}\times n}\underbrace{V}_{n\times d}.$ It is well known that different multiplication sequences can lead to dramatically different numbers of operations required, so the order of matrix multiplications matters, which is indeed the case here. We first perform $V_{1}^{\top}V\in\mathbb{R}^{k_{1}\times d}$ and this cost $O(k_{1}nd)=n^{1+o(1)}$ time. Then we can compute $U_{1}V_{1}^{\top}V$ within $O(nk_{1}d)=n^{1+o(1)}$ time.
This method significantly reduces the computation time of the attention forward from $O(n^{2})$ to almost linear time, $n^{1+o(1)}$. Driven by this technique and analyzing the close forms of the gradients, we extend the acceleration to the gradient of the entire model.
@@ 5.2 Accelerating gradient computation of $T_{i}(X)$
Based on the low-rank approximation method mentioned in Section 5.1, we compute the gradient of $L(X)$ with respect to the intermediate variable $T_{i}(X)$, which denotes the output of the $i$-th transformer layer. This computation is critical as it enables us to calculate gradients for other gradient components because of the chain rule.
Extending to general loss functions.
According to the findings in DSXY [23], the gradient $\frac{\mathrm{d}L(X)}{\mathrm{d}T_{i}(X)}$ can be decomposed into five components, namely $C_{2}(X),C_{4}(X),C_{6}(X),C_{7}(X),C_{8}(X)$, as detailed in Lemma D.1. However, the gradient result presented in DSXY [23] is tailored to a specific loss function, the $\ell_{2}$ loss, limiting its applicability to a narrow range of scenarios. The primary challenge in extending the scope to encompass general loss functions is the absence of a unified analytical framework. Previous analyses [5, 23] are limited to individual, specific loss functions. In this work, we introduce a comprehensive analysis framework (Definition 3.1) and we have demonstrated its applicability to the cross-entropy loss (see Remark 3.2). Consequently, by utilizing this generalized analysis framework, we extend the notation $L(X)$ to include a wide range of general loss functions.
@@
An important step in accelerating the gradient computation for the entire multi-layer transformer model is to accelerate the computation of the gradient on the intermediate variables $T_{i}(X)$. The key challenge is that, to compute the gradient on $T_{i}(X)$, we need to compute the gradient on other components of one transformer layer, such as residual connection, multi-head attention, and causal attention mask (see more details in Section 6). We conduct comprehensive analysis on those components in the transformer layer, and prove that, by using low-rank approximation technique, the computation of gradient $\frac{\mathrm{d}L(X)}{\mathrm{d}T_{i}(X)}$ can be computed in almost linear time $n^{1+o(1)}$ (Lemma 5.1).
A crucial aspect of speeding up gradient computation for the entire multi-layer transformer model involves accelerating the calculation of gradients with respect to the intermediate variables $T_{i}(X)$. The main challenge lies in the fact that computing the gradient of $T_{i}(X)$ requires calculating the gradients for other components within a transformer layer, including the residual connection, multi-head attention, and causal attention mask (see Section 6). We have conducted an extensive analysis of these components within the transformer layer (see Section I, J, and K) and demonstrated that, through the application of low-rank approximation techniques, the gradient $\frac{\mathrm{d}L(X)}{\mathrm{d}T_{i}(X)}$ can be computed in almost linear time $n^{1+o(1)}$ (Lemma 5.1). In particular, we apply the low-rank approximation technique on the five terms $C_{2}(X),C_{4}(X),C_{6}(X),C_{7}(X),C_{8}(X)$ respectively, demonstrating that each term can be computed in almost linear time, $n^{1+o(1)}$, as shown in Section E. Then we aggregate those terms, as described in Section E.6. Since all five terms are $n\times d$ matrices, the summation of these terms remains almost linear in complexity. We then conclude that for any single-layer transformer, the gradient computation with respect to the input can be performed in almost linear time $n^{1+o(1)}$, as stated in Lemma 5.1.
The statement made for a single transformer layer can be readily generalized to any layer within an $m$-layer transformer model. For instance, consider the intermediate variables $T_{i}(X)$ and $T_{i-1}(X)$ (as defined in Definition 3.3), where $T_{i}(X)=(g_{i}\circ\mathsf{Attn}_{i})(T_{i-1}(X))$. Given the gradient $\frac{\mathrm{d}L(X)}{\mathrm{d}T_{i}(X)}$, as established in the previous paragraph, we compute the gradient with respect to $T_{i-1}(X)$, namely $\frac{\mathrm{d}L(X)}{\mathrm{d}T_{i-1}(X)}$, in almost linear time $n^{1+o(1)}$. For a multi-layer transformer model, the above process can be conducted recursively. Thus, we can compute the gradient of the loss function $L(X)$ on any $T_{i}(X)$ in almost linear time $n^{1+o(1)}$.